In 1970 your grandmother put $1,000 into a special trust fund to be paid to a future grandchild (you) 50 years later, in the year 2020. How much will this trust be worth then if it has been earning 9% per year compounded quarterly?
Answer #1

Answer

Formula is:

FV=PV*(1+i)^n

i=effective quarterly interest rate =nominal annual interest rate /number of compounding in the period =9/4=2.25%

n=number of quarters =50*4=200

PV=present value =value deposited in the bank=1000

FV=1000*(1.0225^200)

=85637.3729

the amount is $85637.37